Parker Corporation uses a job-order costing system and applies manufacturing overhead to jobs using a predetermined overhead rate based on direct labor-hours. Last year manufacturing overhead and direct labor-hours were estimated at $80,910 and 27,900 hours, respectively, for the year. In June, Job #461 was completed. Materials costs on the job totaled $4,400 and labor costs totaled $1,836 at $5.40 per hour. At the end of the year, it was determined that the company worked 34,400 direct labor-hours for the year and incurred $89,760 in actual manufacturing overhead costs.

Required:

a. Job #461 contained 100 units. Determine the unit product cost that would appear on the job cost sheet.


b. Determine the underapplied or overapplied overhead for the year.
Answer #1
Predetermined overhead rate = Estimated manufacturing overhead / Estimated direct labor hours = 80910 / 27900 2.9 per direct labor hour
Required a :
Direct labor hours worked on Job #461 = Direct labor cost / Direct labor rate per hour = 1836 / 5.40 340
Job #461
Direct materials 4400
Direct labor cost 1836
Manufacturing overhead applied ( 340 * 2.9 ) 986
Total product cost 7222
(/) Number of units 100
Unit product cost 72.22
Required b :
Total manufacturing overhead applied = Actual direct labor hours worked * Predetermined overhead rate = 34400 * 2.9 99760
Actual manufacturing overhead incurred 89760
As the overhead applied are more than the actual overhead incurred, the overheads are overapplied
Overapplied overhead = Total manufacturing overhead applied - Actual manufacturing overhead incurred = 99760 - 89760   10000